OpenStudy (anonymous):

a stereo system is being installed in a room with a rectangular floor measuring 14 feet by 12 feet and an 8 foot ceiling. the stereo amp is on the floor in one corner of the room. a speaker is a the ceiling in the opposite corner of the room. you must run a wire from the amp to the speaker, and the wire must run along floor or walls. what is the shortest length of wire you can use for the connection?

OpenStudy (anonymous):

is it sqrt(144+196) + 8 = 18.44 +8 =26.44 appx